Question 6 Two particles, each with a charge of magnitude 13.2 nC, are at two of the vertices of an equilateral triangle with edge length 4.7 m. What is the magnitude of the electric field at the third vertex if (a) both charges are positive and (b) one charge is positive and the other is negative? (a) Number (b) Number Units Units
Answer #1

Concept used: we find field due to each charge separately along coordinate axes in component form and add them to find the resultant field,

EA 60* or = 2.689 + 4.6570 -2.68L+4 657 S9 netfeD Ε.et = EAHEB - .315 N/C EA 60* Es 97/ = 2.6891 +4.6570 2. = 2.689 1-4-6570***************************************************************************************************
Check the answer and let me know immediately if you find something wrong... I will rectify the mistakes asap if any